# Preserving Privacy in Large Language Models: A Survey on Current Threats and Solutions

### v1

@@ 7 Discussion and Conclusion
Secondly, we presented an extensive view of the state-of-the-art solutions for preserving privacy in large language models (§4-5). These solutions include approaches that inject privacy directly in the training data like anonymization techniques (§4) or approaches that act on the model (§5). While there are many possible approaches to the problem, no single solution fits all scenarios. We presented solutions for applying privacy mechanisms at several points in the learning pipeline: from data preparation (before training - §4), to private training (during training - §5.1), to private inference (§5.2), and finally after training (§5.4). The main takeaway of our work is that, while various methods address specific needs, the gold standard for privacy-preserving ML is Differential Privacy (DP), which provides a mathematical guarantee of privacy, albeit with some loss in utility and if introduced in the entire learning pipeline (Tramèr et al., 2024; Feng & Tramèr, 2024). DP’s applicability spans data preparation, training, fine-tuning, and inference. In contrast, the privacy guarantee provided by other approaches like Machine Unlearning (§5.4) is only empirical but necessary to implement the "right to be forgotten" of privacy laws like GDPR (Zhang et al., 2024). Future work could focus on establishing a mathematical framework for Machine Unlearning or integrating it with DP, and developing automated systems that allow users to request the deletion of specific information that they want the model to unlearn.
However, despite the extensive application scope of Differential Privacy, it assumes a structure of private data where every bit of language can potentially contain secrets that are not meant to be shared indiscriminately. Differential Privacy cannot directly preserve secrets – information that can be shared only to the correct group of people (Brown et al., 2022b). The capability of LLMs to keep secrets is still little explored, with the development of the first benchmarks to assess the ability of such models to disclose private information only to users who should be allowed to access them (Mireshghallah et al., 2023; Priyanshu et al., 2023; Wang et al., 2024). As LLMs are increasingly integrated as assistants in applications, functioning as agents among groups of users who wish to share certain information while keeping other details private, addressing this issue will become more critical. Future work on Differential Privacy for training, tuning, or conditioning a model should prioritize more recent models with increased parameters and capabilities. While most studies have used smaller models like GPT-2 or BERT, larger models present their own unique challenges, even though recent advances in DP training (e.g., ghost clipping) show promise. Additionally, it would be interesting to see these systems work efficiently in a federated environment with the local model of DP, which is the safest approach but also the least implemented due to computational costs.
@@
In conclusion, this work highlights the critical need to address privacy vulnerabilities in Large Language Models (LLMs) by systematically analyzing various privacy attacks and their profound risks to data confidentiality and user trust. We presented current state-of-the-art solutions such as data anonymization, Differential Privacy, and Machine Unlearning, emphasizing the importance of continuous advancement and integration of these techniques. As LLMs become integral to sensitive domains like healthcare and finance, developing robust privacy safeguards is essential. Future research must enhance these methods, improve the privacy-utility trade-off, and ensure that larger models adhere to stringent privacy standards, ultimately fostering the creation of secure and reliable LLM technologies that uphold user privacy and trust.
8 Acknowledgements
This research is supported by the EU-funded project DataTools4Heart (grant No.101057849).
@@ Appendix A Going Further
As shown throughout this work, the topics discussed here have been subject to extensive research in recent years. In our work, we have shown how privacy-preserving techniques like anonymization or Differential Privacy have been used to build private Language Models. For those readers interested to deepen their knowledge on a single subjects, we present here a list of surveys specific for each single topic:
@@
  • Data Leakage: Jegorova et al. (2022) present a comprehensive survey on privacy leakage at inference time. They cover involuntary leakage and adversarial leakage and also touch on the available defense mechanisms. The main difference with our work is the fact that we are focused on LLMs and that is also reflected on the techniques shown as solutions which are much more recent.
@@
  • Attacks: There are several surveys investigating attacks. Rigaki & García (2020) investigate the privacy vulnerabilities of machine learning in general, while some other works are more focused on a specific topic. For example, Enthoven & Al-Ars (2020); Zhang et al. (2022b); Liu et al. (2022) focus on attacks against federated learning models or Derner et al. (2023); Li et al. (2023a) focus on security risks of LLMs, including possible attacks and mitigations. While some of these works have a partial overlap with our section §3, we mostly focus our investigation on privacy attacks against LLMs and possible solutions.
@@
  • Privacy Preserving Deep Learning: Literature yields many works on the topic of privacy preserving deep learning. Some of them are generic, without focusing on any specific model. This is the case for Boulemtafes et al. (2020); Tanuwidjaja et al. (2020); Tariq et al. (2020); Liu et al. (2021); Tanuwidjaja et al. (2019); Mirshghallah et al. (2020) which all touch on privacy issues in deep learning and possible solutions. Although these works are relevant, they primarily focus on deep learning in general and do not address the recent advancements brought by LLMs. Recently, Yao et al. (2023) provide a systematic review on security and privacy in language models divided into three sections: good applications, bad applications, and security risks and solutions. This work is complementary to our study since it covers a broader range of risks (not just privacy risks) while allocating less space to solutions. Ultimately, works focused on privacy in NLP, such as Sousa & Kern (2022), concentrate on NLP and anonymization, while Guerra-Manzanares et al. (2023) targets medical applications without necessarily focusing on textual data.
@@
  • Machine Unlearning: There are several surveys on the topic of Machine Unlearning: Xu et al. (2023a); Nguyen et al. (2022); Shaik et al. (2024) are general, while Si et al. (2023); Liu et al. (2024) only focus on LLMs and the differences in application.
@@
  • Differential Privacy: With the growing interest in Differential Privacy in recent years, numerous literature reviews have emerged. They all focus on different aspects of Differential Privacy: Prokhorenkov & Cao (2023) are focused on the privacy risks of Differential Privacy, how to define it and measure it. Several works investigated the intersection between Differential Privacy and deep learning like Baraheem & Yao (2022); Zhao et al. (2019b); Ponomareva et al. (2023); Demelius et al. (2023): Chen (2023) focus on generative models in particular; Klymenko et al. (2022) focus on DP in the context of NLP; Garrido et al. (2022) present a particularly interesting work as it focuses on the industry, the difficulties in the adoption and gives some suggestions to close the gap between industry and academia; Ouadrhiri & Abdelhadi (2022); Zhang et al. (2023a); Odera (2023) focus on the applications in federated learning.
@@
  • Federated Learning: Regarding federated learning, several surveys are available with different perspectives. Kairouz et al. (2021b) propose a general survey on federated learning, examining the state of the art and open problems. Hasan (2023); Ma et al. (2022); Truong et al. (2021) give an overview on the privacy and security issues, with Truong et al. (2021) also putting it in relation to GDPR. Lastly Brauneck et al. (2023) give an overview of FL in medical research.
@@ Appendix B Summary Tables
In this section we present summary tables for all the approaches discussed in the text. Table 2 and Table 3 summarize the anonymization techniques respectively presented in §4.1 and §5. Table 4 presents some information about the works that perform a fully private training, discussed in §5.1.1, with information about the model, the datasets used both for pre-training and fine-tuning and the optimization method. Similarly in Table 5 there is an overview of the works that perform private fine-tuning, discussed in §5.1.2. Table 6 summarizes the works found in literature for private inference, which are discussed in §5.2. Lastly, Table 7 presents a brief summary of the Unlearning methods presented in §5.4.
@@
Tramèr & Boneh (2020)
Table 4: Models pre-trained with DP
@@
RoBERTA DP-SGD MNLI, SST-2, QQP, QNLI Yu et al. (2021a)
GPT2 DP-SGD E2E, DART Yu et al. (2021a)
BERT ADAM SST, QQP, TP-UK  [^8] The authors used prompt-tuning and prefix-tuning after obfuscating the embeddings Li et al. (2023d)
T5 ADAM SST, QQP, TP-UK  [^9] The authors used prompt-tuning and prefix-tuning after obfuscating the embeddings Li et al. (2023a)
GPT2 DP-SGD, DP-PPO Reddit TL;DR, IMDb Wu et al. (2024)
Table 5: Models fine-tuned with DP
@@
Table 7: Machine Unlearning Methods

### v2

@@ 8 Conclusion
Secondly, we presented an extensive view of the state-of-the-art solutions for preserving privacy in large language models (§4-5). These solutions include approaches that inject privacy directly into the training data, like anonymization techniques (§4) or approaches that act on the model (§5). While there are many possible approaches to the problem, no single solution fits all scenarios. We presented solutions for applying privacy mechanisms at several points in the learning pipeline: from data preparation (before training - §4), to private training (during training - §5.1), to private inference (§5.2-5.5), and finally after training (§5.4). We refer the reader to section 7 for a thorough discussion on current limitations and future directions for preserving privacy in large language models.
@@
This work highlights the critical need to address privacy vulnerabilities in Large Language Models (LLMs) by systematically analyzing various privacy attacks and their profound risks to data confidentiality and user trust. We presented current state-of-the-art solutions such as data anonymization, Differential Privacy, and Machine Unlearning, emphasizing the importance of continuously advancing and integrating these techniques. As LLMs become integral to sensitive domains like healthcare and finance, developing robust privacy safeguards is essential. Future research must enhance these methods, improve the privacy-utility trade-off, and ensure that larger models adhere to stringent privacy standards, ultimately fostering the creation of secure and reliable LLM technologies that uphold user privacy and trust.
Acknowledgements
The authors gratefully acknowledge the support of the EU-funded project DataTools4Heart (grant No. 101057849), which provided crucial funding and resources for this research. This work has also been carried out in the scope of Project ECS 0000024 Rome Technopole – CUP B83C22002820006, NRP Mission 4 Component 2 Investment 1.5, funded by the European Union – NextGenerationEU.
We also extend our sincere thanks to the reviewers and the action editor of TMLR for their invaluable feedback and constructive comments. Their insights have significantly enhanced the quality and clarity of this paper.
@@ Appendix A Going Further
As shown throughout this work, the topics discussed here have been subject to extensive research in recent years. In our work, we have shown how privacy-preserving techniques like anonymization or Differential Privacy have been used to build private Language Models. For those readers interested in deepening their knowledge on a single subject, we present here a list of surveys specific to each single topic:
@@
  • Data Leakage: Jegorova et al. (2021) presents a comprehensive survey on privacy leakage at inference time. They cover involuntary leakage and adversarial leakage and also touch on the available defence mechanisms. The main difference with our work is that we are focused on LLMs, which is also reflected in the techniques shown as solutions, which are much more recent.
@@
  • Attacks: There are several surveys investigating attacks. Rigaki & García (2020) investigate the privacy vulnerabilities of machine learning in general, while some other works are more focused on a specific topic. For example, Enthoven & Al-Ars (2021); Zhang et al. (2022b); Liu et al. (2022) focus on attacks against federated learning models or Derner et al. (2023); Li et al. (2023b) focus on security risks of LLMs, including possible attacks and mitigations. While some of these works have partial overlap with our section §3, we mostly focus our investigation on privacy attacks against LLMs and possible solutions.
@@
  • Privacy Preserving Deep Learning: Literature yields many works on the topic of privacy-preserving deep learning. Some of them are generic, without focusing on any specific model. This is the case for Boulemtafes et al. (2020); Tanuwidjaja et al. (2020); Tariq et al. (2020); Liu et al. (2021); Tanuwidjaja et al. (2019); Mirshghallah et al. (2020) which all touch on privacy issues in deep learning and possible solutions. Although these works are relevant, they primarily focus on deep learning in general and do not address the recent advancements brought by LLMs. Recently, Yao et al. (2024b) provided a systematic review on security and privacy in language models divided into three sections: good applications, bad applications, and security risks and solutions. This work is complementary to our study since it covers a broader range of risks (not just privacy risks) while allocating less space to solutions. Ultimately, works focused on privacy in NLP, such as Sousa & Kern (2023), concentrate on NLP and anonymization, while Guerra-Manzanares et al. (2023) targets medical applications without necessarily focusing on textual data.
@@
  • Machine Unlearning: There are several surveys on the topic of Machine Unlearning: Xu et al. (2023a); Nguyen et al. (2022); Shaik et al. (2024) are general, while Si et al. (2023); Liu et al. (2024a) only focus on LLMs and the differences in application.
@@
  • Differential Privacy: With the growing interest in Differential Privacy in recent years, numerous literature reviews have emerged. They all focus on different aspects of Differential Privacy: Prokhorenkov & Cao (2023) is focused on the privacy risks of Differential Privacy, how to define it and measure it. Several works investigated the intersection between Differential Privacy and deep learning like Baraheem & Yao (2022); Zhao et al. (2019b); Ponomareva et al. (2023); Demelius et al. (2023): Chen et al. (2024a) focus on generative models in particular; Klymenko et al. (2022) focus on DP in the context of NLP; Garrido et al. (2023) present a particularly interesting work as it focuses on the industry, the difficulties in the adoption and gives some suggestions to close the gap between industry and academia; Ouadrhiri & Abdelhadi (2022); Zhang et al. (2023a); Odera (2023) focus on the applications in federated learning.
@@
  • Federated Learning: Regarding federated learning, several surveys are available with different perspectives. Kairouz et al. (2021b) propose a general survey on federated learning, examining the state of the art and open problems. Hasan (2023); Ma et al. (2022); Truong et al. (2021) give an overview of the privacy and security issues, with Truong et al. (2021) also putting it in relation to GDPR. Lastly, Brauneck et al. (2023) gives an overview of FL in medical research.
@@ Appendix B Summary Tables
In this section, we present summary tables for all the approaches discussed in the text. Table 2 and Table 3 summarize the anonymization techniques respectively presented in §4.1 and §5. Table 4 presents some information about the works that perform a fully private training, discussed in §5.1.1, with information about the model, the datasets used both for pre-training and fine-tuning and the optimization method. Similarly, in Table 5, there is an overview of the works that perform private fine-tuning, discussed in §5.1.2. Table 6 summarizes the works found in the literature for private inference, which are discussed in §5.2. Lastly, Table 7 presents a brief summary of the Unlearning methods presented in §5.4.
@@
Tramer & Boneh (2021)
@@
RoBERTA DP-SGD MNLI, SST-2, QQP, QNLI Yu et al. (2022)
GPT2 DP-SGD E2E, DART Yu et al. (2022)
BERT ADAM SST, QQP, TP-UK  [^8] The authors used prompt-tuning and prefix-tuning after obfuscating the embeddings Li et al. (2023e)
T5 ADAM SST, QQP, TP-UK  [^9] The authors used prompt-tuning and prefix-tuning after obfuscating the embeddings Li et al. (2023b)
GPT2 DP-SGD, DP-PPO Reddit TL;DR, IMDb Wu et al. (2024a)
@@
MEMIT
With a similar approach with respect to ROME,
scale the editing on multiple examples.
The discussion is focused on factuality.
Meng et al. (2022)
DEPN
This method assumes that private information is
encoded in specific neurons: their weights are set
to zero to block information flow.
Wu et al. (2023)
PAE
Model editing to break the association between
a user identity and its PII. The edit method is
effective both in batch and in sequential editing.
Venditti et al. (2024)
@@
Appendix C Discussion on Federated Learning
McMahan et al. (2018) introduce DP-FedAvg, a variation of the FedAVG algorithm by McMahan et al. (2016), incorporating the moments accountant (Abadi et al., 2016) to calculate the privacy budget used during training. Typically, FedAVG performs part of the training locally on a subset of clients selected based on specific criteria. The clients then send their updated gradients to the server, which averages them, updates the global model, and sends the updated model back to the clients. In DP-FedAvg, gradients are clipped locally but not obfuscated. The central server adds noise after receiving and aggregating the gradients to prevent inversion, aligning the algorithm with central DP since local clipping alone does not meet the Local DP criteria. This approach was tested by training an LSTM for the next token prediction on a large dataset of Reddit posts described by Al-Rfou et al. (2016), using various noise levels and clipping strategies. The model’s performance was evaluated using AccuracyTop1, the probability that the predicted word is correct. The study demonstrated that, with sufficient data, it is possible to train a model with privacy guarantees without significantly sacrificing performance, provided the noise added is kept within reasonable limits.
Currently, large language models (LLMs) represent the state of the art due to their superior power, though they are much larger than the small LSTMs seen in previous studies. Wang et al. (2024b) focus on using LLMs to enhance the privacy/utility tradeoff of local LMs in differentially private federated learning. Similar to Yu et al. (2024), which samples pre-training data based on its similarity to fine-tuning data, they propose a distribution matching algorithm to sample public data close to the private data distribution, significantly improving sample efficiency. They utilize DP-FTRL (Kairouz et al., 2021a), an online optimization algorithm combining FTL (follow the leader) and regularization, adding noise during local updates. The local LM is either a single-layer LSTM (Hochreiter & Schmidhuber, 1997) or Transformer (Vaswani et al., 2017) due to computational limits on edge devices, while the LLM used is LaMDA (Thoppilan et al., 2022). Their tests yielded several notable results. First, using a public tokenizer with a public vocabulary instead of a private one helps prevent privacy leakage. Second, pre-training local models on a public dataset (C4 dataset) for over a week of single GPU time improved next-token prediction accuracy (tested on the private StackOverflow dev set). They also built a distillation corpus by saving the top-k logits with k nonzero entries from the teacher LLM’s next token predictions as a silver label dataset. Pre-training on this distillation corpus showed accuracy comparable to pre-training with the original dataset, even with just 1% distillation coverage. Furthermore, their new distribution matching technique allows using only 0.08% of the pre-training dataset to achieve performance equal to 1% distillation coverage, greatly enhancing sample efficiency.
Xu et al. (2023b) present and analyze twenty Gboard LMs trained for Next Word Prediction using DP-FTRL, asserting that all future Gboard models will have DP guarantees. In their DP-FTRL implementation, model updates are clipped, and noise is added during global model updates, ensuring Central Differential Privacy. They also use adaptive clipping (Andrew et al., 2021), which adjusts the clip norm each round by privately estimating the model’s norm at a targeted quantile. However, adaptive clipping showed no significant improvement and was less robust in experiments with large report goals, sometimes causing catastrophic failure or no progress in the first 1000 rounds. Despite this, adaptive clipping can reduce hyperparameter tuning when the privacy budget permits. Public pre-training on C4, as explained by Wang et al. (2024b), reduced the rounds needed to reach target utility by  1000 under the same noise multiplier, enhancing privacy guarantees. They also combined DP-FTRL with Secure Aggregation (Bonawitz et al., 2017), a cryptographic protocol ensuring the central server only accesses aggregated updates, not individual client updates. While SecAgg aids DP, it introduces significant computational slowdown and requires different system configurations for large report goals. The techniques for training language models with differential privacy in Gboard have produced models with stronger privacy guarantees than some established standards, such as those used by the US Census Bureau. In Europe, these differentially private neural network models outperform older N-gram models in utility. For English in the US and Portuguese in Brazil, the new models are on par with or slightly better than their non-private counterparts.
C.1 Federated Learning with Local DP
Youn et al. (2023) introduce RQM (Randomized Quantization Mechanism), a novel approach to enforce DP in federated settings with improved communication efficiency, used alongside Distributed SGD on a CNN for image classification on EMNIST. RQM enhances DP-SGD by eliminating the need for explicit noise addition; instead, RQM is performed after gradient clipping. In RQM, gradient updates are encoded in quantization levels, which are then randomly subsampled. A randomized rounding procedure maps continuous gradient values to the nearest subsampled quantization level. RQM falls under Renyi Differential Privacy (Mironov, 2017), a relaxation of the original differential privacy definition, placing the method under Local DP since RQM is performed on the clients. Tests show that RQM outperforms the Poisson Binomial Mechanism (Chen et al., 2022b) (a traditional noise injection method) and noise-free clipped SGD, achieving better accuracy with stronger privacy guarantees.